OpenStudy (anonymous):

what is the rule for fractions is the following problem? I am so lost in this class already... 1/2x + 1/3 = 5/6 give me tips and not answer?

OpenStudy (anonymous):

in*

jimthompson5910 (jim_thompson5910):

in the first fraction, is the denominator 2 or 2x?

OpenStudy (anonymous):

the denominator is 2 but the fraction has an x next to it

jimthompson5910 (jim_thompson5910):

so the equation is \[\large \frac{1}{2}x + \frac{1}{3} = \frac{5}{6}\] right?

OpenStudy (anonymous):

yes =)

jimthompson5910 (jim_thompson5910):

what are the denominators

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

2,3,6

jimthompson5910 (jim_thompson5910):

what is the LCM of those numbers

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

so the LCD is 6

jimthompson5910 (jim_thompson5910):

what you can do from here is multiply each fraction by the LCD and that will eliminate the fractions

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

12x+18=5/6

jimthompson5910 (jim_thompson5910):

no, that's incorrect

jimthompson5910 (jim_thompson5910):

what is 6 times 1/2

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

6*1/2 = (6/1)*(1/2) 6*1/2 = (6*1)/(1*2) 6*1/2 = 6/2 6*1/2 = 3

jimthompson5910 (jim_thompson5910):

so 6 times 1/2 = 3

jimthompson5910 (jim_thompson5910):

what is 6 times 1/3

OpenStudy (anonymous):

its 8

jimthompson5910 (jim_thompson5910):

hint: 6*1/3 = (6/1)*(1/3)

jimthompson5910 (jim_thompson5910):

it's not 8

OpenStudy (anonymous):

I am so bad at this class . please pray for me. its 2

jimthompson5910 (jim_thompson5910):

you'll be fine, just keep up the hard work and you'll get better

jimthompson5910 (jim_thompson5910):

good, 6 times 1/3 = 2

jimthompson5910 (jim_thompson5910):

what is 6 times 5/6

OpenStudy (anonymous):

1 Remainder 1?

jimthompson5910 (jim_thompson5910):

6 times 5/6 = (6/1)*(5/6) = ??

OpenStudy (anonymous):

5?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

So after multiplying EVERY fraction by the LCD 6, we go from \[\large \frac{1}{2}x + \frac{1}{3} = \frac{5}{6}\] to \[\large 3x + 2 = 5\]

jimthompson5910 (jim_thompson5910):

so you can see how multiplying every term by the LCD eliminates all the fractions

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

I got it from here. thank you

jimthompson5910 (jim_thompson5910):

ok that's great, yw

OpenStudy (anonymous):

hello? I have another problem if you want to help? I am recording in my brain.

jimthompson5910 (jim_thompson5910):

sure what's the question